# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v2

@@ 1 Introduction
  • Low-rank compression framework. We introduce a novel regularization and incorporation method to modify a class of low-rank training frameworks that yields low-rank compressed neural networks, achieving a more than $10\times$ reduction in both memory footprint and compute cost, while maintaining clean accuracy and adversarial robustness on par with full-rank baselines.
@@
Beyond these core contributions, our approach is model- and data-agnostic, can be integrated seamlessly with existing adversarial defenses, e.g., adversarial training goodfellow2014explaining , and never requires assembling full-rank weight matrices – the last point guaranteeing a low memory footprint during training and inference. Moreover, by connecting to dynamical low-rank integration schemes and enabling convergence analysis via gradient flow, we offer new theoretical and algorithmic insights. Finally, the use of interpretable spectral metrics enhances the trustworthiness and analyzability of the compressed models.
@@
We consider a neural network $f$ as a concatenation of $L$ layers $z^{\ell+1}=\sigma^{\ell}(W^{\ell}z^{\ell})$ with matrix valued [^1] We provide an extension to tensor-valued layers, e.g. in CNNs, in Section 5.1 parameters $W^{\ell}\in\mathbb{R}^{n\times n}$, layer input $z^{\ell}\in\mathbb{R}^{n\times b}$ and element-wise nonlinear activation $\sigma^{\ell}$. For simplicity of notation, we do not consider biases, but they are included in the numerical experiments in Section˜6. The data $X$ constitutes the input to the first layer, i.e. $z^{0}=X$. We assume that the layer activations $\sigma^{\ell}$ are Lipschitz continuous, which is the case for all popular activations savostianova2023robustlowranktrainingapproximate . The network is trained on a loss function $\mathcal{L}$ which we assume to be locally bounded with a Lipschitz continuous gradient. Throughout this work, we call a network in the standard format a "baseline" network.
Low-rank Compression: The compression the network for training and inference is typically facilitated by approximating the layer weight matrices by a low-rank factorization $W^{\ell}=U^{\ell}S^{\ell}V^{\ell,\top}$ with $U^{\ell},V^{\ell}\in\mathbb{R}^{n\times r}$ and $S^{\ell}\in\mathbb{R}^{r\times r}$, where $r$ is the rank of the factorization. In this work, we generally assume that $U^{\ell},V^{\ell}$ are orthonormal matrices at all times during training and inference. This assumption deviates from standard low-rank training approaches hu2021lora , however recent literature provides methods that are able to fulfill this assumption approximately zhang2023adalora and even exactly steffen2022low ; schotthöfer2024geolorageometricintegrationparameter . If $r\ll n$, the low-rank factorization with a storage and matrix-vector computation cost cost of $\mathcal{O}(2nr+r^{2})$ is computationally more efficient than the standard matrix format $W$ with a computational cost of $\mathcal{O}(n^{2})$.
@@
Figure 1: The singular values $\varsigma_{i}(W)$ of sequential layer 7 in VGG16 for baseline training, unregularized dynamical low-rank training (DLRT), and RobustDLRT with our condition number regularizer $\mathcal{R}$ with $\beta=0.075$. The matrix $W$ is formed as the first-mode unfolding of the convolutional tensor. Conditioning of the regularized low-rank layer is significantly improved compared to the non-regularized low-rank and baseline layer.2
@@
The sensitivity of a low-rank factorized network can be readily deducted from Equation˜1 by leveraging orthonormality of $U$ and $V$, i.e., $\kappa(USV^{\top})=\kappa(S)$. This fact allows us to only consider the $r\times r$ coefficient matrix $S$ to control the sensitivity of the network. The condition number $\kappa(S)$ can be determined via a singular value decomposition (SVD) of $S$, which is computationally feasible when $r\ll n$.
Adversarial robustness-aware low-rank training: Enhancing the adversarial robustness of the network during low-rank training thus boils down to controlling the conditioning of $S$, which is a non-trivial task. In fact, the dynamics of the singular spectrum of $S$ of adaptive low-rank training schemes as DLRT steffen2022low become more ill-conditioned than the baseline during training, even if $S$ is always full rank. In Figure˜1, we observe that the singular values $\varsigma$ of a rank 64 factorization of a network layer compressed with DLRT range from $\varsigma_{r=1}=2.7785$ to $\varsigma_{r=64}=0.8210$ yielding a condition number of $\kappa(S)=3.3844$. In comparison, the baseline network has singular values ranging from $\varsigma_{r=1}=1.8627$ to $\varsigma_{r=128}=0.9445$ yielding a lower condition number of $\kappa(S)=1.9722$. As a result, an $\ell^{2}$-FGSM attack with strength $\epsilon=0.3$, reduces the accuracy of the baseline network to $54.96\%$, while the low-rank network’s accuracy drops to $43.39\%$, see Table˜2.
@@ 3 Related work
Low-rank compression is a prominent approach for reducing the memory and computational cost of deep networks by constraining weights to lie in low-rank subspaces. Early methods used post-hoc matrix denton2014exploiting and tensor decompositions (lebedev2015speeding, ), while more recent approaches integrate low-rank constraints during training for improved efficiency and generalization.
@@
Dynamical Low-Rank Training (DLRT) steffen2022low constrains network weights to evolve on a low-rank manifold throughout training, allowing substantial reductions in memory and FLOPs without requiring full-rank weight storage. The method has been extended to tensor-valued neural network layers zangrando2023rank , and federated learning schotthöfer2024federateddynamicallowranktraining . Pufferfish (wang2021pufferfish, ) restricts parameter updates to random low-dimensional subspaces, while intrinsic dimension methods (aghajanyan2020intrinsic, ) argue that many tasks can be learned in such subspaces. GaLore (zhao2024galore, ) reduces memory cost by projecting gradients onto low-rank subspaces.
@@
In contrast, low-rank fine-tuning methods like LoRA (hu2021lora, ) inject trainable low-rank updates into a frozen pre-trained model, enabling efficient adaptation with few parameters. Extensions such as GeoLoRA schotthöfer2024geolorageometricintegrationparameter , AdaLoRA (zhang2023adalora, ), DyLoRA (valipour2022dylora, ), and DoRA (Mao2024DoRAEP, ) incorporate rank adaptation or structured updates, improving performance over static rank baselines. However, these fine-tuning methods do not reduce the cost of the full training and inference, thus are not applicable to address the need of promoting computational efficiency.
@@
Pruning is another well studied approach to reduce the number of parameters of a trained neural network jian2022pruningadversariallyrobustneural ; li2022pruningimprovecertifiedrobustness ; NEURIPS2020_e3a72c79 ; pmlr-v188-zhao22a ; chen2022linearitygraftingrelaxedneuron ; jordao2021effect by either sparsifying weight matrices or layer output channels of a network. Typically sparsity pruning is performed after training a fully parametrized neural network and thus only reduces memeory and compute load during inference, while treating training as offline cost.
@@
Improving adversarial robustness with orthogonal layers has been a recently studied topic in the literature anil2019sorting ; bansal2018can ; xie2017all ; cisse17_parseval ; savostianova2023robustlowranktrainingapproximate . Many of these methods can be classified as either a soft approach, where orthogonality is imposed weakly via a regularizer, or a hard approach where orthogonality is explicitly enforced in training.
@@
Examples of soft approaches include the soft orthogonal (SO) regularizer xie2017all , double soft orthogonal regularizer bansal2018can , mutual coherance regularizer bansal2018can , and spectral normalization miyato2018spectral . These regularization-based approaches have several advantages; namely, they are more flexible to many problems/architectures and are amenable to transfer learning scenarios (since pertained models are admissible in the optimization space). However, influencing the spectrum weakly via regularization cannot enforce rigorous and explicit bounds on the spectrum.
@@
Many hard approaches strongly enforce orthogonality/well-conditioned constraints by training on a chosen manifold using Riemannian optimization methods li2020efficientriemannianoptimizationstiefel ; Projection_like_Retractions ; savostianova2023robustlowranktrainingapproximate . A hard approach built for low-rank training is given in savostianova2023robustlowranktrainingapproximate ; this method clamps the extremes of the spectrum to improve the condition number during training. The clamping gives a hard estimate on the range of the spectrum which enables a direct integration of the low-rank equations of motion with standard learning rates. However, this method requires a careful selection of the rank $r$, which is viewed as a hyperparameter in savostianova2023robustlowranktrainingapproximate . If $r$ is chosen incorrectly, the clamping of the spectrum, a hard-thresholding technique, acts as a strong regularizer which could affect the validation metrics of the network.
@@
Our regularization method detailed below falls neatly into a soft approach and our proposed regularizer can be seen as an extension of the soft orthogonality (SO) regularizer xie2017all to well-conditioned matrices in the low-rank setting. As noted in bansal2018can , the SO regularizer only works well when the input matrix is of size $m\times n$ wish $m\leq n$. However, in our case we avoid this issue since the regularizer is applied to the square matrix $S$; an extension to convolutional layers is discussed in Section˜5.1. In the context of low-rank training, the soft approach enables rank-adaptivity of the method.
@@ 4 Improving conditioning via regularization
${\mathcal{R}}(S)=\|S^{\top}\!S-\alpha_{S}^{2}I\|\qquad\text{where}\qquad\alpha_{S}^{2}=\frac{1}{r}\|S\|^{2}$ (2)
@@
The regularizer $\mathcal{R}$} can be viewed as an extension of the soft orthogonal regularizer xie2017all ; bansal2018can where we penalize the distance of $S^{\top}\!S$ to the well-conditioned matrix $\alpha_{S}^{2}I$. Here $\alpha_{S}$ is chosen such that $\|S\|=\|\alpha_{S}I\|$. Moreover, $\mathcal{R}$} is also a scaled standard deviation of the singular values $\{\varsigma_{i}(S)^{2}\}_{i=1}^{r}$, i.e.,
@@
See Appendix˜C for the proof. The gradient computation consists only of $r\times r$ matrix multiplications and an Frobenius norm evaluation, and is thus computationally efficient for $r\ll m$. Further, its closed form enables a straight-forward integration into existing optimizers as Adam or SGD applied to $S$.
@@
See Appendix˜C for the proof. Thus, if $\varsigma_{r}(S)$ is not too small, we can use ${\mathcal{R}}(S)$ as a good measure for the conditioning of $S$. Note that the singular value truncation used in rank-adaptive methods ensures that $\varsigma_{r}(S)$ is always sufficiently large. Figure˜2a) and b) show the dynamics of ${\mathcal{R}}(S(t))$ and $\kappa(S(t))$ during low-rank regularized training; we see that $\kappa(S(t))$ decays as ${\mathcal{R}}(S(t))$ decays, validating ˜2.
@@
When $U,V$ are not orthonormal, e.g., in simultaneous gradient descent training (LoRA), Moreover, the smallest nrn-r singular values of $USV^{\top}$ are often zero-valued, and thus the bound of Equation˜4 is not useful. Table˜1 shows that the clean accuracy and adversarial accuracy of regularized LoRA is significantly lower than standard or regularized training with orthonormal $U,V$.
@@
5 A rank-adaptive, adversarial robustness increasing dynamical low-rank training scheme
@@
To facilitate the discussion, we define $\widetilde{\mathcal{L}}=\mathcal{L}+\beta\mathcal{R}$ as the regularized loss function of the training process with regularization parameter $\beta>0$. To construct the method we consider the (stochastic) gradient descent-based update of a single weight matrix $W_{t+1}=W_{t+1}-\lambda\nabla_{W}\widetilde{\mathcal{L}}$ for minimizing $\widetilde{\mathcal{L}}$ with step size $\lambda>0$. The corresponding continuous time gradient flow reads $\dot{W}(t)=-\nabla_{W}\widetilde{\mathcal{L}}(W(t))$, which is a high-dimensional dynamical system with a steady state solution. We draw from established dynamical low-rank approximation (DLRA) methods, which were initially proposed for matrix-valued equations KochLubich07 . DLRA was recently extended to neural network training steffen2022low ; zangrando2023rank ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; kusch2025augmentedbackwardcorrectedprojectorsplitting ; Hnatiuk to formulate a consistent gradient flow evolution for the low-rank factors $U$, $S$, and $V$.
@@
Below we summarize the proposed method, called RobustDLRT, along with the changes induced by adding our robustness regularizer. A single iteration of RobustDLRT is specified in Algorithm˜1.
@@
to double the rank of the low-rank representation and subsequently creates orthonormal bases $^{\widehat{U}},{\widehat{V}}$. Since ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$, $\nabla_{U}{\mathcal{R}}(USV^{\top})=\nabla_{V}{\mathcal{R}}(USV^{\top})=0$; hence $\nabla_{U}\widetilde{\mathcal{L}}=\nabla_{U}\mathcal{L}$ and $\nabla_{V}\widetilde{\mathcal{L}}=\nabla_{V}\mathcal{L}$ are used in (6). The span of $^{\widehat{U}}$ contains $U^{t}$, which is needed to ensure of the loss does not increase during augmentation, and a first-order approximation of $\mathrm{span}(U^{t+1})$ using the exact gradient flow for $U$, see (schotthöfer2024federateddynamicallowranktraining, , Theorem 2) for details. Geometrically, the latent space
@@
We update the latent coefficients $^{\widehat{S}}$ via a Galerkin projection of the training dynamics onto the latent space $\mathcal{S}$. The latent coefficients $^{\widehat{S}}$ are updated by integrating the projected gradient flow $\dot{{\widehat{S}}}=-{\widehat{U}}^{\top}\nabla_{W}\widetilde{\mathcal{L}}{\widehat{V}}=-\nabla_{\widehat{S}}\widetilde{\mathcal{L}}$ using stochastic gradient descent or an other suitable optimizer for a number of $s_{*}$ local iterations, i.e.
@@
The computational cost of RobustDLRT is asymptotically the same as LoRA, since the reconstruction of the full weight matrix $W$ is never required. The orthonormalization accounts for $\mathcal{O}(nr^{2})$, the regularizer $\mathcal{R}$ for $\mathcal{O}(r^{3})$, and the SVD for $\mathcal{O}(r^{3})$ floating point operations. When using multiple coefficient update steps $s_{*}>1$, the amortized cost is indeed lower than that of LoRA, since only the gradient with respect to $^{\widehat{S}}$ is required in most updates. While the regularizer may be applied to full-rank baseline models, its $\mathcal{O}(n^{3})$ computational scaling significantly increases training costs.

### v3

@@ 1 Introduction
  • Low-rank compression framework. We introduce a novel regularization and integration method to modify a class of low-rank training methods that yields low-rank compressed neural networks, achieving a more than $10\times$ reduction in both memory footprint and compute cost, while maintaining clean accuracy and adversarial robustness on par with full-rank baselines.
@@
Beyond these core contributions, our approach is model- and data-agnostic, can be integrated seamlessly with existing adversarial defenses, e.g., adversarial training goodfellow2014explaining , and never requires assembling full-rank weight matrices—the last point guaranteeing a low memory footprint during training and inference. Moreover, by connecting to dynamical low-rank integration schemes and enabling convergence analysis via gradient flow, we offer new theoretical and algorithmic insights. Finally, the use of interpretable spectral metrics enhances the trustworthiness and analyzability of the compressed models.
@@
We consider a neural network $f$ as a concatenation of $L$ layers $z^{\ell+1}=\sigma^{\ell}(W^{\ell}z^{\ell})$ with matrix valued [^1] We provide an extension to tensor-valued layers, e.g. in CNNs, in Section 5.1 parameters $W^{\ell}\in\mathbb{R}^{n\times n}$, layer input $z^{\ell}\in\mathbb{R}^{n\times b}$ and element-wise nonlinear activation $\sigma^{\ell}$. For simplicity of notation, we do not consider biases, but they are included for the numerical experiments in Section˜6. The data $X$ constitutes the input to the first layer, i.e. $z^{0}=X$. We assume that the layer activations $\sigma^{\ell}$ are Lipschitz continuous, which is the case for all popular activations savostianova2023robustlowranktrainingapproximate . The network is trained on a loss function $\mathcal{L}$ which we assume to be locally bounded with a Lipschitz continuous gradient. Throughout this work, we call a network in the standard format a “baseline” network.
Low-rank Compression: The compression the network for training and inference is typically facilitated by approximating the layer weight matrices by a low-rank factorization $W^{\ell}=U^{\ell}S^{\ell}V^{\ell,\top}$ with $U^{\ell},V^{\ell}\in\mathbb{R}^{n\times r}$ and $S^{\ell}\in\mathbb{R}^{r\times r}$, where $r\leq n$ is the rank of the factorization. In this work, we generally assume that $U^{\ell},V^{\ell}$ are orthonormal matrices at all times during training and inference. This assumption deviates from standard low-rank training approaches hu2021lora , however recent literature provides methods that are able to fulfill this assumption approximately zhang2023adalora and even exactly steffen2022low ; schotthöfer2024geolorageometricintegrationparameter . If $r\ll n$, the low-rank factorization with a storage and matrix-vector computation cost cost of $\mathcal{O}(2nr+r^{2})$ is computationally more efficient than the standard matrix format $W$ with a computational cost of $\mathcal{O}(n^{2})$.
@@
Figure 1: The singular values $\varsigma_{i}(W)$ of sequential layer 7 in VGG16 for baseline training, unregularized dynamical low-rank training, and RobustDLRT with our condition number regularizer $\mathcal{R}$ with $\beta=0.075$ (see Section˜5). The matrix $W$ is formed as the first-mode unfolding of the convolutional tensor. Conditioning of the regularized low-rank layer is significantly improved compared to the non-regularized low-rank and baseline layer.2
@@
The sensitivity of a low-rank factorized network can be readily deducted from Equation˜1 by leveraging orthonormality of $U$ and $V$, i.e., $\kappa(USV^{\top})=\kappa(S)$. Thus, we only consider the $r\times r$ coefficient matrix $S$ to control the sensitivity of the network. The condition number $\kappa(S)$ can be determined via a singular value decomposition (SVD) of $S$, which is computationally feasible when $r\ll n$.
Adversarial robustness-aware low-rank training: Enhancing the adversarial robustness of the network during low-rank training thus boils down to controlling the conditioning of $S$, which is a non-trivial task. Moreover, the dynamics of the singular spectrum of $S$ of adaptive low-rank training schemes as Dynamical Low-Rank Training (DLRT) steffen2022low become more ill-conditioned than the baseline during training, even if $S$ is always full rank. In Figure˜1, we observe that the singular values $\varsigma$ of a rank 64 factorization of a network layer compressed with DLRT range from $\varsigma_{r=1}=2.7785$ to $\varsigma_{r=64}=0.8210$ yielding a condition number of $\kappa(S)=3.3844$. In comparison, the baseline network has singular values ranging from $\varsigma_{r=1}=1.8627$ to $\varsigma_{r=128}=0.9445$ yielding a lower condition number of $\kappa(S)=1.9722$. As a result, an $\ell^{2}$-FGSM attack with strength $\epsilon=0.3$, reduces the accuracy of the baseline network to $54.96\%$, while the accuracy of the low-rank network drops to $43.39\%$, see Table˜2.
@@ 3 Related work
Low-rank compression is a prominent approach for reducing the memory and computational cost of deep networks by constraining weights to lie in low-rank subspaces. Early methods used post-hoc matrix denton2014exploiting and tensor decompositions (lebedev2015speeding, ), while more recent approaches enforce low-rank constraints during training for improved efficiency and generalization.
@@
Dynamical Low-Rank Training steffen2022low constrains network weights to evolve on a low-rank manifold throughout training, allowing substantial reductions in memory and FLOPs without requiring full-rank weight storage. The method has been extended to tensor-valued neural network layers zangrando2023rank , and federated learning schotthöfer2024federateddynamicallowranktraining . Pufferfish (wang2021pufferfish, ) restricts parameter updates to random low-dimensional subspaces, while intrinsic dimension methods (aghajanyan2020intrinsic, ) argue that many tasks can be learned in such subspaces. GaLore (zhao2024galore, ) reduces memory cost by projecting gradients onto low-rank subspaces.
@@
In contrast, low-rank fine-tuning methods like low-rank adaptation (LoRA) (hu2021lora, ) inject trainable low-rank updates into a frozen pre-trained model, enabling efficient adaptation with few parameters. Extensions such as GeoLoRA schotthöfer2024geolorageometricintegrationparameter , AdaLoRA (zhang2023adalora, ), DyLoRA (valipour2022dylora, ), and DoRA (Mao2024DoRAEP, ) incorporate rank adaptation or structured updates, improving performance over static rank baselines. However, these fine-tuning methods do not reduce the cost of the full training and inference, thus are not applicable to address the need of promoting computational efficiency.
@@
Pruning is another well studied approach to reduce the number of parameters of a trained neural network jian2022pruningadversariallyrobustneural ; li2022pruningimprovecertifiedrobustness ; NEURIPS2020_e3a72c79 ; pmlr-v188-zhao22a ; chen2022linearitygraftingrelaxedneuron ; jordao2021effect by either sparsifying weight matrices or layer output channels of a network. Typically sparsity pruning is performed after training a fully parametrized neural network and thus only reduces memory and compute load during inference, while treating training as an offline cost.
@@
Improving adversarial robustness with orthogonal layers has been a recently studied topic in the literature anil2019sorting ; bansal2018can ; xie2017all ; cisse17_parseval ; savostianova2023robustlowranktrainingapproximate . Many of these methods can be classified as either a soft approach, where orthogonality is imposed weakly via a regularizer, or a hard approach, where orthogonality is explicitly enforced in training.
@@
Examples of soft approaches include the soft orthogonal (SO) regularizer xie2017all , double soft orthogonal regularizer bansal2018can , mutual coherence regularizer bansal2018can , and spectral normalization miyato2018spectral . These regularization-based approaches have several advantages; namely, they are more flexible to many problems/architectures and are amenable to transfer learning scenarios (since pertained models are admissible in the optimization space). However, influencing the spectrum weakly via regularization cannot enforce rigorous and explicit bounds on the spectrum.
@@
Many hard approaches strongly enforce orthogonality/well-conditioned constraints by training on a chosen manifold using Riemannian optimization methods li2020efficientriemannianoptimizationstiefel ; Projection_like_Retractions ; savostianova2023robustlowranktrainingapproximate . A hard approach built for low-rank training is given in savostianova2023robustlowranktrainingapproximate ; this method clamps the extremes of the spectrum to improve the condition number during training. The clamping gives a hard estimate on the range of the spectrum which enables a direct integration of the low-rank equations of motion with reasonable learning rates. However, this method requires a careful selection of the rank $r$, which is viewed as a hyperparameter in savostianova2023robustlowranktrainingapproximate . If $r$ is chosen incorrectly, the clamping of the spectrum, a hard-thresholding technique, acts as a strong regularizer which could affect the validation metrics of the network.
@@
Our regularization method detailed below falls neatly into a soft approach and our proposed regularizer can be seen as an extension of the soft orthogonality (SO) regularizer xie2017all to well-conditioned matrices in the low-rank setting. As noted in bansal2018can , the SO regularizer only works well when the input matrix is of size $m\times n$ with $m\leq n$. However, we avoid this issue since the regularizer is applied to the square $r\times r$ matrix $S$; an extension to convolutional layers is discussed in Section˜5.1. In the context of low-rank training, the soft approach enables rank-adaptivity of the method.
@@ 4 Improving conditioning via regularization
${\mathcal{R}}(S)=\|S^{\top}\!S-\alpha_{S}^{2}I\|,\qquad\text{where}\qquad\alpha_{S}^{2}=\frac{1}{r}\|S\|^{2}$ (2)
@@
The regularizer $\mathcal{R}$} can be viewed as an extension of the soft orthogonal regularizer xie2017all ; bansal2018can where we penalize the distance of $S^{\top}\!S$ to the well-conditioned matrix $\alpha_{S}^{2}I$. Here $\alpha_{S}$ is chosen such that $\|S\|=\|\alpha_{S}I\|$. Moreover, $\mathcal{R}$} is also a scaled standard deviation of the squared singular values $\{\varsigma_{i}(S)^{2}\}_{i=1}^{r}$, i.e.,
@@
See Appendix˜C for the proof. The gradient computation consists only of $r\times r$ matrix multiplications and a Frobenius norm evaluation. Thus $\nabla$$\mathcal{R}$} is computationally efficient for $r\ll m$. Further, its closed form enables a straight-forward integration into existing optimizers such as Adam or SGD applied to $S$.
@@
See Appendix˜C for the proof. Thus, if $\varsigma_{r}(S)$ is not too small, we can use ${\mathcal{R}}(S)$ as a good measure for the conditioning of $S$. Note that the singular value truncation used in rank-adaptive methods ensures that $\varsigma_{r}(S)$ is always sufficiently large. Figures˜2(a) and 2(b) show the dynamics of ${\mathcal{R}}(S(t))$ and $\kappa(S(t))$ during low-rank regularized training; we see that $\kappa(S(t))$ decays as ${\mathcal{R}}(S(t))$ decays, validating ˜2.
@@
When $U,V$ are not orthonormal, e.g., in simultaneous gradient descent training (LoRA), the smallest nrn-r singular values of $USV^{\top}$ are often zero-valued; thus, the bound of Equation˜4 is not useful. Table˜1 shows that the clean accuracy and adversarial accuracy of regularized LoRA is significantly lower than standard or regularized training with orthonormal $U,V$.
@@
5 A rank-adaptive and adversarial robustness increasing dynamical low-rank training scheme
@@
To facilitate the discussion, we define $\widetilde{\mathcal{L}}=\mathcal{L}+\beta\mathcal{R}$ as the regularized loss function of the training process with regularization parameter $\beta>0$. To construct the method we consider the (stochastic) gradient descent-based update of a single weight matrix $W_{t+1}=W_{t+1}-\lambda\nabla_{W}\widetilde{\mathcal{L}}$ for minimizing $\widetilde{\mathcal{L}}$ with step size $\lambda>0$. The corresponding continuous time gradient flow reads $\dot{W}(t)=-\nabla_{W}\widetilde{\mathcal{L}}(W(t))$, which is a high-dimensional dynamical system with a steady state solution. We draw from established dynamical low-rank approximation (DLRA) methods, which were initially proposed for matrix-valued dynamical systems KochLubich07 . DLRA was recently extended to neural network training steffen2022low ; zangrando2023rank ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; kusch2025augmentedbackwardcorrectedprojectorsplitting ; Hnatiuk to formulate a consistent gradient flow evolution for the low-rank factors $U$, $S$, and $V$.
@@
Below we list the method of schotthöfer2024federateddynamicallowranktraining with the changes introduced by adding our robustness regularizer. We call the resulting scheme RobustDLRT, and a single iteration of RobustDLRT is specified in Algorithm˜1.
@@
to double the rank of the low-rank representation and subsequently creates orthonormal bases $^{\widehat{U}},{\widehat{V}}$. Here $\texttt{orth}(A)$ denotes an orthonormal basis for the range of $A$ and $\mid$ denotes horizontal concatenation of matrices. Since ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$, $\nabla_{U}{\mathcal{R}}(USV^{\top})=\nabla_{V}{\mathcal{R}}(USV^{\top})=0$; hence $\nabla_{U}\widetilde{\mathcal{L}}=\nabla_{U}\mathcal{L}$ and $\nabla_{V}\widetilde{\mathcal{L}}=\nabla_{V}\mathcal{L}$ are used in (6). The span of $^{\widehat{U}}$ contains $U^{t}$, which is needed to ensure of the loss does not increase during augmentation, and a first-order approximation of $\mathrm{span}(U^{t+1})$ using the exact gradient flow for $U$, see (schotthöfer2024federateddynamicallowranktraining, , Theorem 2) for details. Geometrically, the latent space
@@
We update the latent coefficients $^{\widehat{S}}$ via a Galerkin projection of the training dynamics onto the latent space $\mathcal{S}$. The latent coefficients $^{\widehat{S}}$ are updated by integrating the projected gradient flow $\dot{{\widehat{S}}}=-{\widehat{U}}^{\top}\nabla_{W}\widetilde{\mathcal{L}}{\widehat{V}}=-\nabla_{\widehat{S}}\widetilde{\mathcal{L}}$ using stochastic gradient descent or an other suitable optimizer for a number of $s_{*}$ local iterations, i.e.,
@@
The computational cost of RobustDLRT is asymptotically the same as LoRA, since the reconstruction of the full weight matrix $W$ is never required. The orthonormalization, computation of the regularizer $\mathcal{R}$, and the SVD for accounts for $\mathcal{O}(nr^{2})$, $\mathcal{O}(r^{3})$, $\mathcal{O}(r^{3})$ floating point operations, respectively. When using multiple coefficient update steps $s_{*}>1$, the amortized cost is lower than that of LoRA, since only the gradient with respect to $^{\widehat{S}}$ is required in most updates. While the regularizer may be applied to full-rank baseline models, its $\mathcal{O}(n^{3})$ computational scaling significantly increases training costs.